# Pure C*-algebras

### v1

@@ 3. Variants of comparison
We first recall the definition of $m$-comparison as considered by Toms and Winter in [TW09, Lemma 6.1], and later formalized by Ortega, Perera and Rørdam in [OPR12, Definition 2.8].
@@
Let $m\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$. The lowest possible value for $m$-comparison is attained when $m=0$, which is commonly referred to as saying that $S$ is almost unperforated.
We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does.
@@
(1) L. Robert showed in [Rob11, Theorem 1.3] that, if $A$ is a $C^{*}$-algebra with nuclear dimension at most $m$, then $A$ has $m$-comparison.
@@
Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$; see [APRT24, Notation 5.2] and also [APT18, Definition 5.6.1]. To recall the definition of locally bounded comparison amplitude, we need the notion of a scale in a $\mathrm{Cu}$-semigroup.
3.3Scales.
Let $S$ be a $\mathrm{Cu}$-semigroup. A scale in $S$ is a subset $\Sigma\subseteq S$ such that
@@
    is downward hereditary: if $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$,
@@
    is closed under suprema of increasing sequences, and
@@
(See [APT20c, Definition 4.1]). We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and
@@
We recall below the definition of locally bounded comparison amplitude, from [APRT24]. Useful reformulations of this condition will be given in Lemma 3.5.
Definition 3.4 ([APRT24, Definition 6.1]).
A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has locally bounded comparison amplitude, in short (LBCA), if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
@@
A $C^{*}$-algebra will be said to have locally bounded comparison amplitude if $(\operatorname{Cu}(A),\Sigma_{A})$ has (LBCA).
@@
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup. The following are equivalent:
  1. (i)
    $(S,\Sigma)$ has locally bounded comparison amplitude.
  2. (ii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that
    $\widehat{x}\leq\gamma\widehat{y^{\prime}},\quad y^{\prime}\ll y\quad\text{ implies }\quad Nx\leq Ny,\qquad\text{for all }x,y^{\prime},y\in\Sigma^{(d)}.$
  3. (iii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $M=M(\gamma,d)$ such that:
    $\widehat{x}\leq\gamma\widehat{y}\quad\text{ implies }\quad(M+1)x\leq My,\qquad\text{for all }x,y\in\Sigma^{(d)}.$
@@
It is clear that (i) implies (ii).
Let us show that (ii) implies (iii). Fix $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$. Choose $\gamma^{\prime}$ and $n\in{\mathbb{N}}$ such that
$\gamma<\gamma^{\prime},\quad\frac{n+1}{n}\gamma^{\prime}<1.$
Let $N=N(\tfrac{n+1}{n}\gamma^{\prime},d(n+1))$ as provided by (ii). We claim that $M=Nn$ has the desired properties to verify (iii) (for $\gamma$ and $d$).
To this end, let $x,y\in\Sigma^{(d)}$ satisfy $\widehat{x}\leq\gamma\widehat{y}$. Let $x^{\prime}\in S$ be such that $x^{\prime}\ll x$. By Proposition 2.3 (i) applied to $x^{\prime}\ll x$ and $1<\frac{\gamma^{\prime}}{\gamma}$, we have that $\widehat{x^{\prime}}\ll\gamma^{\prime}\widehat{y}$. Hence, there exists $y^{\prime}\ll y$ such that $\widehat{x^{\prime}}\leq\gamma^{\prime}\widehat{y^{\prime}}$. Then
$(n+1)\widehat{x^{\prime}}\leq(n+1)\gamma^{\prime}\widehat{y^{\prime}}=\left(\frac{n+1}{n}\gamma^{\prime}\right)n\widehat{y^{\prime}}.$
Since $(n+1)x^{\prime},ny^{\prime}$ and $ny$ belong to $\Sigma^{((n+1)d)}$, it follows from the choice of $N$ that $N(n+1)x^{\prime}\leq Nny$. Passing to the supremum over all $x^{\prime}\ll x$ we get that $N(n+1)x\leq Nny$, and thus $(M+1)x\leq My$.
Finally, note that if elements $x$ and $y$ in a partially ordered semigroup satisfy $(M+1)x\leq My$ for some $M\in{\mathbb{N}}$, then there exists $N\in{\mathbb{N}}$ such that $(n+1)x\leq ny$ for all $n\geq N$, and in particular $x\leq_{N}y$; see [APT18, Proposition 5.2.13]. This shows that (iii) implies (i). ∎
The following lemma is implied by [APRT24, Proposition 10.3].
Lemma 3.6.
Any $C^{*}$-algebra with $m$-comparison has locally bounded comparison amplitude.
Proof.
In order to apply [APRT24, Proposition 10.3] it is required that the Cuntz semigroup of a $C^{*}$-algebra satisfies (O5), (O6), and the so-called Edward’s condition. The first two are always satisfied, as mentioned in the preliminaries. As for the third condition, this was also shown to hold in general in [APRT21, Theorem 5.3]. ∎
Theorem 3.7.
If a $C^{*}$-algebra has finite nuclear dimension, then its Cuntz semigroup has locally bounded comparison amplitude.
Proof.
Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$. As observed in Remarks 3.2, $\operatorname{Cu}(A)$ has $m$-comparison, hence the result follows from Lemma 3.6. ∎
4. Variants of divisibility
We first recall the definition of $n$-almost divisibility for $\mathrm{Cu}$-semigroups, which appeared in [RT17, Section 2.3]. Other related notions of divisibility will be also useful in our analysis. These have sprung into prominence due to their connection with important structural properties of $C^{*}$-algebras.
Definition 4.1.
Let $k\in{\mathbb{N}}$, and $n\in{\mathbb{N}}\cup\{\infty\}$. We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is $(k,n)$-divisible if, for each $x^{\prime}\in S$ with $x^{\prime}\ll x$, there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(n+1)y.$
If all elements are $(k,n)$-divisible, when $n=\infty$, then we say that $S$ is $(k,\omega)$-divisible.
If all elements are $(k,n)$-divisible for all $k$, we then say that $S$ is $n$-almost divisible. In case $S$ is 00-almost divisible, then we just say that $S$ is almost divisible.
A $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is.
Remarks 4.2.
(1) It was proved in [TV24, Theorem 9.1] that, if $A$ is a nowhere scattered $C^{*}$-algebra of real rank zero, then $A$ is almost divisible. In fact, its Cuntz semigroup $\operatorname{Cu}(A)$ is weakly divisible: for any $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6]. In the simple case, and for the Murray-von Neumann semigroup, this was partly obtained in [PR04, Proposition 5.3].
(2) If a $\mathrm{Cu}$-semigroup $S$ is almost unperforated, then it is almost divisible precisely when, given $x\in S$ and $k$, there is $y\in S$ such that $ky\leq x\leq(k+1)y$. This was proved in [APT18, Proposition 7.3.7].
(3) The concept of $(k,\omega)$-divisibility is connected to the Global Glimm Property which, in the context of this paper, will be analysed with more detail in Section 6. It was proved in [TV23, Theorem 3.6] that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible.
Inspired by the results in [RT17, Section 6], we introduce the following more general divisibility conditions in terms of comparison by functionals.
Definition 4.3.
We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is
  • (i)
@@ 4. Variants of divisibility
    functionally almost divisible, or that $x$ has (FAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, every $k\in{\mathbb{N}}_{>0}$, and every $\varepsilon>0$, there exists $y\in S$ such that
  • (ii)
    lower functionally almost divisible, or that $x$ has (LFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
  • (iii)
    upper functionally almost divisible, or that $x$ has (UFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ and $\nu>0$ such that
    $y\leq x,\quad(k+\nu)\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)y.$
A $\mathrm{Cu}$-semigroup $S$ is said to be (upper, lower) functionally almost divisible if every element in $S$ is (upper, lower) functionally almost divisible.
We begin our study of functional divisibility by providing a number of technical results relating each of the notions above. These results will be applied to $C^{*}$-algebras with $n$-almost divisibility at the end of the section, where we will show that any $n$-almost divisible $C^{*}$-algebra satisfies all three divisibility notions; see Theorem 4.7.
Lemma 4.4.
Let $S$ be a $\mathrm{Cu}$-semigroup.
  1. (i)
    $S$ is functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\tfrac{1}{k}\widehat{x}.$
  2. (ii)
    $S$ is lower functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+\tfrac{1}{k})\widehat{y}.$
Proof.
We will only prove (i) since the proof of (ii) is similar. Moreover the forward implications are clear. To prove the backwards implication, let $x^{\prime},x\in S$ satisfy $x^{\prime}\ll x$, let $k\in{\mathbb{N}}_{>0}$, and let $\varepsilon>0$. Choose $n\in{\mathbb{N}}_{>0}$ such that $\tfrac{1}{kn}\leq\varepsilon$. Applying the assumption for $x^{\prime},x$ and $kn$, we obtain $z\in S$ such that
$knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}.$
Set $y:=nz$. Then
$ky=knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}=k\widehat{y}+\frac{1}{kn}\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x},$
as desired. ∎
Lemma 4.5.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ is $(2,\omega)$-divisible and is functionally almost divisible. Then $S$ is lower functionally almost divisible.
Proof.
Let $k\in{\mathbb{N}}$, and let $x^{\prime},x\in S$ be such that $x^{\prime}\ll x$. It suffices to prove that there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
For $k=1$ use $y:=x$. Thus, we may from now on assume that $k\geq 2$.

### v3

@@ 3. Variants of comparison
In this section we deal with various comparison properties. In particular, we recall the definition of $m$-comparison (first considered in [TW09, Lemma 6.1] and later formalized in [OPR12, Definition 2.8]) and clarify the equivalence between 00-comparison (also known as almost unperforation) with the notion of strict comparison for $C^{*}$-algebras. We also recall the concept introduced in [APRT24, Definition 6.1], which in this paper is renamed as controlled comparison.
@@
Let $m\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$.
We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does. A $\mathrm{Cu}$-semigroup or $C^{*}$-algebra is said to be almost unperforated if it has 00-comparison.
@@
(1) Robert showed in [Rob11, Theorem 1.3] that a $C^{*}$-algebra has $m$-comparison whenever it has nuclear dimension at most $m$.
@@
The notion of strict comparison was first considered in [Bla88]. For $\mathrm{Cu}$-semigroups, the concept was considered in [ERS11, Section 6].
Definition 3.3.
A $\mathrm{Cu}$-semigroup $S$ has strict comparison if for all $x,y\in S$ satisfying $x\leq\infty y$ and $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ with $\lambda(y)=1$, one has $x\leq y$.
A $C^{*}$-algebra $A$ has strict comparison of positive elements by quasitraces (specifically, $[0,\infty]$-valued, lower semicontinuous $2$-quasitraces) if its Cuntz semigroup does.
3.4Strict comparison and almost unperforation.
We take the opportunity here to clarify the equivalence between the concepts of strict comparison and almost unperforation. Let $S$ be a $\mathrm{Cu}$-semigroup, and let $x,y\in S$. The following conditions relating the comparison of $x,y$ and $\widehat{x},\widehat{y}$ were considered in [APT18]:
@@
    We have $x<_{s}y$, that is, there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$.
@@
    We have $\widehat{x}<_{s}\widehat{y}$, that is, there exists $\varepsilon\in(0,1)$ such that $\widehat{x}\leq(1-\varepsilon)\widehat{y}$.
  3. (iii)
    We have $x\leq\infty y$ and $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ with $\lambda(y)=1$.
  4. (iv)
    We have $x^{\prime}<_{s}y$ for every $x^{\prime}\in S$ with $x^{\prime}\ll x$.
It was shown in [APT18, Theorem 5.2.18] that the implications (i)$\Rightarrow$(ii)$\Rightarrow$(iii)$\Rightarrow$(iv) always hold. Using this, it follows that the following conditions are equivalent for $S$. (This was first obtained in [ERS11, Proposition 6.2]; see also [APT18, Proposition 5.2.20].)
  1. (i’)
    $S$ is almost unperforated.
  2. (ii’)
    If $x,y\in S$ satisfy $\widehat{x}\leq(1-\varepsilon)\widehat{y}$ for some $\varepsilon\in(0,1)$, then $x\leq y$.
  3. (iii’)
    $S$ has strict comparison.
Now, let $A$ be a $C^{*}$-algebra. There is a natural correspondence between functionals on $\operatorname{Cu}(A)$ and quasitraces on $A$. This is defined by sending a quasitrace $\tau\colon(A\otimes\mathcal{K})_{+}\to[0,\infty]$ to the functional $d_{\tau}\colon\operatorname{Cu}(A)\to[0,\infty]$ given by
$d_{\tau}([a])=\lim_{n\to\infty}\tau(a^{1/n})$
for $a\in(A\otimes\mathcal{K})_{+}$; see, for example, [ERS11, Proposition 4.2].
Using this, it follows that $A$ has strict comparison if and only if for all $a,b\in(A\otimes\mathcal{K})_{+}$ such that $a$ belongs to the closed ideal generated by $b$, and such that $d_{\tau}([a])<d_{\tau}([b])$ for every quasitrace $\tau$ with $d_{\tau}([b])=1$, we have $a\precsim b$.
It also follows from the above considerations that a $C^{*}$-algebra $A$ has strict comparison (of positive elements by quasitraces) if and only if it is almost unperforated.
Some of the confusion surrounding the notion of strict comparison arises since it is applied in a more restrictive setting, for example by considering comparison of projections (instead of positive elements), or by considering comparison by traces (instead of quasitraces); see [NR16, Section 3] for some clarifications.
Another source of confusion is that strict comparison is often used in the context of simple $C^{*}$-algebras, in which case every element is automatically contained in the closed ideal generated by a nonzero element. Therefore, a simple, unital $C^{*}$-algebra $A$ has strict comparison if and only if for all nonzero elements $a,b\in(A\otimes\mathcal{K})_{+}$ such that $d_{\tau}([a])<d_{\tau}([b])$ for every normalized quasitrace $\tau$, we have $a\precsim b$.
For the definition below, we need to recall the notion of a scale in a $\mathrm{Cu}$-semigroup.
3.5Scales.
Let $S$ be a $\mathrm{Cu}$-semigroup. Following [APT20c, Definition 4.1], a scale in $S$ is a subset $\Sigma\subseteq S$ satisfying the following conditions:
  1. (i)
    $\Sigma$ is downward hereditary: If $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$.
  2. (ii)
    $\Sigma$ is closed under suprema of increasing sequences.
@@
We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and for $d\geq 1$ as
@@
Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$. The following was termed ‘locally bounded comparison amplitude’ in [APRT24, Definition 6.1].
Definition 3.6.
A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has controlled comparison if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
@@
We say that a $C^{*}$-algebra $A$ has controlled comparison if $(\operatorname{Cu}(A),\Sigma_{A})$ does.
The next result follows easily from [APRT24, Proposition 10.3].
Proposition 3.7.
If a $C^{*}$-algebra has $m$-comparison for some $m\in{\mathbb{N}}$, then it has controlled comparison.
Proof.
In order to apply [APRT24, Proposition 10.3] it is required that the Cuntz semigroup of a $C^{*}$-algebra satisfies (O5), (O6), and the so-called Edward’s condition. The first two are always satisfied, as mentioned in the preliminaries, and the third condition was shown to hold in [APRT21, Theorem 5.3]. ∎
Corollary 3.8.
Every $C^{*}$-algebra with finite nuclear dimension has controlled comparison.
Proof.
Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$. Then $A$ has $m$-comparison by [Rob11, Theorem 1.3], hence the result follows from Proposition 3.7. ∎
4. Variants of divisibility
In this section we introduce the concept of functional divisibility and show it is a natural generalization of $n$-almost divisibility (as introduced in [Win12, Definition 3.5] for simple, unital $C^{*}$-algebras, and later refined in [RT17, Section 2.3] for general $C^{*}$-algebras). We prove that functional divisibility is automatic for nowhere scattered $C^{*}$-algebras of stable rank one, and also for simple, unital, non-elementary $C^{*}$-algebras with a unique normalized quasitrace.
Definition 4.1.
Let $n\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ is $n$-almost divisible if for every $x^{\prime}\ll x$ in $S$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(n+1)y.$
We say that a $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is. A $\mathrm{Cu}$-semigroup or $C^{*}$-algebra is said to be almost divisible if it is 00-almost divisible. In other words, if for all $x^{\prime}\ll x$ in $S$ and $k\in{\mathbb{N}}$ with $k\geq 1$ there exists $y\in S$ such that $ky\leq x$, and $x^{\prime}\leq(k+1)y$.
4.2Connections with earlier divisibility notions.
In parts of the literature, ‘almost divisible’ is defined as the slightly stronger condition that for every $x\in S$ and $k\in{\mathbb{N}}$ with $k\geq 1$ there exists $y\in S$ such that
$ky\leq x\leq(k+1)y.$
This version of ‘almost divisibility’ implies the one from Definition 4.1, and for almost unperforated $\mathrm{Cu}$-semigroups both versions agree by [APT18, Proposition 7.3.7]. The advantage of the version from Definition 4.1 is that it enjoys better permanence properties; see [APT18, Remark 7.3.5].
Almost divisibility is automatic in certain cases. For example, it was shown in [TV24a, Theorem 9.1] that every nowhere scattered $C^{*}$-algebra $A$ of real rank zero is almost divisible. In fact, $\operatorname{Cu}(A)$ is even ‘weakly divisible’ in the sense that for every $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6]. In the simple case, and for the Murray-von Neumann semigroup, this was obtained in [PR04, Proposition 5.3].
Inspired by the results in [RT17, Section 6] and also in [Win12], we introduce the following more general divisibility condition in terms of comparison by functionals.
Definition 4.3.
We say that a $\mathrm{Cu}$-semigroup $S$ is functionally divisible if the following conditions are satisfied:
  • (L)
    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
  • (U)
    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon\in(0,1)$, there exists $z\in S$ such that
    $z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
We say that a $C^{*}$-algebra $A$ is functionally divisible if $\operatorname{Cu}(A)$ is.
Remark 4.4.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$. While Definition 4.1 captures notions of divisibility of $x$ in $S$, one may also consider the divisibility of $\widehat{x}$ in $\operatorname{Lsc}(F(S))$. The straightforward concept would be to require that for every $x\in S$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that $k\widehat{y}=\widehat{x}$. Taking also the way-below relation into account in order to obtain better permanence properties, one is led to the following notion:
  • (D)
    For every $x\in S$, every $f\in\operatorname{Lsc}(F(S))$ with $f\ll\widehat{x}$, and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that
    $f\leq k\widehat{y}\leq\widehat{x}.$
Using Proposition 2.3(i), we see that a function $f\in\operatorname{Lsc}(F(S))$ satisfies $f\ll\widehat{x}$ if and only if there exists $x^{\prime}\in S$ with $x^{\prime}\ll x$ and $\varepsilon>0$ such that $f\leq(1-\varepsilon)\widehat{x^{\prime}}$. Thus, condition (D) holds if and only if for every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
$(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}\leq\widehat{x}.$
Conditions (L) and (U) in Definition 4.3 are natural strengthenings of this, where one of the inequalities on functionals is upgraded to an inequality in $S$. The ‘L’ stands for ‘lower’, since in this condition the element $ky$ is smaller than $x$, while the ‘U’ stands for ‘upper’ since here $ky$ is larger than $x^{\prime}$.
The next result shows that it suffices to consider $k=2$ in order to verify condition (L) in Definition 4.3.
Proposition 4.5.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5). Then the following statements are equivalent:
  1. (i)
    $S$ satisfies condition (L) in Definition 4.3.
  2. (ii)
    For every $x^{\prime}\ll x$ in $S$ and $k\in{\mathbb{N}}$, there exists $y\in S$ such that
  3. (iii)
    For every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$ there exists $y\in S$ such that
    $2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y}.$
Proof.
Assuming (i), let us verify (ii). For $k=0$, we use $y=x$. For $k\geq 1$, we apply (i) with $\varepsilon>0$ small enough such that $(1-\varepsilon)^{-1}k\leq k+1$. Similarly, to see that (ii) implies (iii), for a given $\varepsilon\in(0,1)$ apply (ii) with $2k$ where $k$ is large enough such that $1-\varepsilon\leq(1+\tfrac{1}{2k})^{-1}$.
Next, assuming that (iii) holds, let us verify (i). First, by induction over $m$, we verify that the following statement holds:
[…]
    For every $x^{\prime}\ll x$ in $S$ and $\varepsilon\in(0,1)$ there exists $y\in S$ such that
    $2^{m}y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2^{m}\widehat{y}.$
[…]
$x^{\prime}\ll x^{\prime\prime}\ll x,\,\,\,\text{ and }\,\,\,(1-\delta)^{3}=1-\varepsilon.$
[…]
$2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime\prime}}\leq 2^{m}\widehat{w}.$
By Proposition 2.3(i), we have $(1-\delta)^{2}\widehat{x^{\prime}}\ll(1-\delta)\widehat{x^{\prime\prime}}$ in $\operatorname{Lsc}(F(S))$, hence also $(1-\delta)^{2}\widehat{x^{\prime}}\ll 2^{m}\widehat{w}$. This allows us to choose $w^{\prime}\in S$ such that
$w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,(1-\delta)^{2}\widehat{x^{\prime}}\leq 2^{m}\widehat{w^{\prime}}.$
Next, applying (iii) for $w^{\prime}\ll w$ and $\delta$, we obtain $y\in S$ such that
$2y\leq w,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{w^{\prime}}\leq 2\widehat{y}.$
Then
$2^{m+1}y\leq 2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}=(1-\delta)^{3}\widehat{x^{\prime}}\leq(1-\delta)2^{m}\widehat{w^{\prime}}\leq 2^{m+1}\widehat{y},$
[…]
Now, to prove (i), let $x^{\prime}\ll x$ in $S$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$. Pick $x^{\prime\prime}\in S$ and $\delta>0$ such that
$x^{\prime}\ll x^{\prime\prime}\ll x,\,\,\,\text{ and }\,\,\,(1-\delta)^{2}=1-\varepsilon.$
Using that dyadic rationals are dense in $\mathbb{R}$, choose $n,m\geq 1$ such that
$(1-\delta)\frac{1}{k}\leq\frac{n}{2^{m}}\leq\frac{1}{k}.$
Then
$(1-\delta)2^{m}\leq kn\leq 2^{m}.$
[…]
$2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime}}\leq 2^{m}\widehat{w}.$
Set $y:=nw$. Then
$ky=k(nw)\leq 2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}=(1-\delta)^{2}\widehat{x^{\prime}}\leq(1-\delta)2^{m}\widehat{w}\leq kn\widehat{w}=k\widehat{y},$
which shows that $y$ has the desired properties. ∎
Remark 4.6.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5). It is not clear if there exists a characterization of condition (U) in Definition 4.3 analogous to Proposition 4.5. The reason is that condition (U) has the additional assumption that the dividing element is dominated by $x$, which is automatic in condition (L).
On the other hand, for $m\geq 1$ let us consider the following condition:
  • ($U_{2^{m}}$)
    For every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
    $y\leq x,\quad(1-\varepsilon)2^{m}\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq 2^{m}y.$
An argument as in the proof of Proposition 4.5 shows that if $S$ satisfies ($U_{2}$), then it satisfies ($U_{2^{m}}$) for all $m$. (See also Remark 5.6.)
To show that $n$-almost divisible Cuntz semigroups are functionally divisible (Proposition 4.9), we first need some preparatory technical results. We will use the concept of $(2,\omega)$-divisibility (see also Section 6). Given $k\in{\mathbb{N}}$ with $k\geq 2$, a $\mathrm{Cu}$-semigroup $S$ is said to be $(k,\omega)$-divisible if for all $x^{\prime}\ll x$ in $S$ there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq\infty y.$
If $S$ is $(k,\omega)$-divisible for some $k\geq 2$, then it is $(2,\omega)$-divisible. The converse also holds by [TV23, Lemma 3.4]. Thus, a $\mathrm{Cu}$-semigroup is $(2,\omega)$-divisible if and only if it is $(k,\omega)$-divisible for all $k\geq 2$.
Condition (L’) in the result below was considered in [RT17, Theorem 6.1].
@@ 4. Variants of divisibility
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5)-(O7). Then $S$ satisfies condition (L) in Definition 4.3 if and only if $S$ is $(2,\omega)$-divisible and satisfies
  1. (L’)
    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
@@
To show the forward implication, assume that $S$ satisfies
  • (L)
@@
    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}\setminus\{0\}$, and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
To verify that $S$ is $(2,\omega)$-divisible, let $x^{\prime}\ll x$ in $S$. Applying (L) with the given $x^{\prime},x$, and also $k=2$ and $\varepsilon=\tfrac{1}{2}$, we obtain $y\in S$ such that
$2y\leq x,\,\,\,\text{ and }\,\,\,\frac{1}{2}\widehat{x^{\prime}}\leq 2\widehat{y}.$
The latter condition implies that $\widehat{x^{\prime}}\leq 4\widehat{y}$, and hence $x^{\prime}\leq\infty y$ (using, for example, [APRT22, Lemma 6.6]). This shows that $S$ is $(2,\omega)$-divisible.
To verify (L’), let $x^{\prime}\ll x$ in $S$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and $\varepsilon>0$. Pick $\delta\in(0,1)$ such that $(1-\delta)^{-1}=1+\varepsilon$. Applying (L) with the given $x^{\prime},x$ and $k$, and also $\delta$, we obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime}}\leq k\widehat{y}.$
Using that $ky\leq x$, we get
$\widehat{x^{\prime}}\leq(1-\delta)^{-1}k\widehat{y}=(1+\varepsilon)k\widehat{y}\leq k\widehat{y}+\varepsilon\widehat{x},$
as desired.
To show the backward implication, assume that $S$ is $(2,\omega)$-divisible and satisfies (L’). We verify condition (ii) of Proposition 4.5. Let $x^{\prime}\ll x$ in $S$, and let $k\in{\mathbb{N}}$. We need to find $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+1)\widehat{y}.$
For $k=0$ and $k=1$ use $y:=x$. Thus, from now on we may assume that $k\geq 2$.